# SculptDrug : A Spatial Condition-Aware Bayesian Flow Model for Structure-based Drug Design

### v1

@@ 1 Introduction
Drug discovery is the process of identifying potential therapeutic molecules to treat diseases, playing a crucial role in improving health and addressing unmet medical needs [1]. Yet, this process is highly complex and time-consuming. Structure-Based Drug Design (SBDD) has emerged as a powerful approach to streamline this process [2], leveraging the structural information of biological targets to rationally design molecules with improved specificity and efficacy [3].
The efficacy of SBDD lies in its ability to leverage the lock-and-key principle of protein-ligand interactions [4], providing a framework for designing molecules that precisely target biological systems, as shown in Figure 1. Specifically, the protein surface functions as a unique “lock”, while the ligand serves as the complementary “key”. This structural compatibility ensures precise protein-ligand interactions, highlighting the effectiveness of SBDD in developing targeted and efficient drugs.
@@
Traditionally, SBDD relies on virtual screening, which aims to select ligands from an immense chemical space estimated to contain approximately $10^{60}$ potential molecules [5]. virtual screening focuses on predefined compound libraries based on molecular properties and structural features [6], which cover only a small fraction of the space and significantly limit ligand discovery [5]. Recent deep generative models [7, 8, 9, 10] enable ligand generation beyond predefined libraries, supporting broader exploration of chemical space [11].
@@
Diffusion models and Bayesian Flow Networks (BFNs) [10] are powerful tools for molecular generation, but often struggle with (3) Spatial Modeling Fidelity. Unlike images with fixed spatial layouts, molecules rely on flexible interatomic distances to define structure. Noise added to atomic coordinates can distort these distances, causing atoms to fall outside interaction thresholds and be excluded from local computations, ultimately compromising chemical validity.
@@ 2 Related Work
Ligand generation has evolved from 1D representations like SMILES [12, 1], to 2D molecular graphs [13, 14], and more recently to 3D structure-based approaches. While 1D/2D methods capture some chemical constraints, they lack the spatial detail essential for modeling protein-ligand interactions. With the rise of deep learning in 3D modeling, structure-aware ligand generation has become a central research focus. Existing methods can be broadly categorized into: (1) Voxel-based and (2) Euclidean space-based models.
Voxel-based methods discretize 3D space into regular grids and model ligand density distributions within this voxelized space. Representative works such as LIGAN [15] and VoxBind [16] generate ligands aligned to protein binding pockets using voxel-based representations. However, these methods suffer from high computational cost at fine resolutions and limited spatial precision due to the loss of structural detail during voxelization.
Euclidean space-based methods generate atomic types and positions directly in continuous 3D coordinates, typically representing proteins and ligands as point clouds or graphs. Early models, including GraphBP [17], Pocket2Mol [18], and FLAG [19], adopt autoregressive strategies at either the atom or fragment level. SurfGen [20] and ResGen [21] incorporate protein context through surface features or residue-level graphs. However, the autoregressive nature of these methods can lead to error accumulation and limited global awareness.
Diffusion-based approaches, such as DIFFBP [22], TARGETDIFF [23], and DIFFSBDD [24], leverage 3D-equivariant denoising to iteratively refine atom placements, improving spatial consistency and structural integrity. Extensions like D3FG [25] and DECOMPDIFF [26] introduce fragment priors and scaffold decomposition to enhance diversity and controllability. Nonetheless, modeling discrete atom types alongside continuous coordinates remains challenging. The introduction of Bayesian Flow Networks adds a new dimension to molecular generation  [27]. MOLCRAFT [28] generates ligands in continuous parameter spaces. However, their limited ability to perceive protein spatial geometry constrains performance in structure-guided ligand design.
@@ 4.1 Boundary Awareness Block
Inspired by SurfGen [20] and SurfPro [29], we extract the surface structure of the binding pocket by selecting residues within 10 Å of any ligand atom. The solvent-excluded surface is computed using MSMS [30], which outputs a triangular mesh of surface vertices and faces. We retain only the inward-facing vertices and their adjacent edges that are spatially close to the ligand. Each vertex is annotated with geometric and biochemical features: the shape index is computed from local curvature, while hydrophobicity, polarity, and electrostatic charge are assigned based on the nearest residue. The resulting surface graph aligns with the formal definition of $\mathcal{S}$.
@@ 4.2 Hierarchical Encoder
While atomic point clouds provide fine-grained local details, they often fail to capture higher-order structural patterns and are sensitive to noise. To address this, we first generate virtual atoms by clustering protein atoms, producing a coarse-grained representation that aggregates within-cluster features and highlights global structure. Ligands then interact with these virtual atoms to integrate global context, while fine-grained interactions are captured via multi-type edges that model detailed biochemical cues. To support both levels of interaction, we adopt equivariant attention mechanisms in the global and local modules, drawing inspiration from the attention design in DecompDiff [26].
@@
To obtain a coarse-grained protein representation, we apply k-means++ [31] clustering to protein atoms based on 3D coordinates, generating virtual atoms at cluster centroids. Let $C$ denote the atoms in a cluster. The virtual atom’s position is defined as:
@@
Note: Evina >> 0 are considered invalid and are represented by “–”. Baseline ligand results are from CBGbench [32].
@@
We designed fine-grained edge connection rules grounded in established domain knowledge, specifically inspired by the affinity calculations in AutoDock Vina [33]. This approach ensures that our modeling of molecular interactions is both accurate and biologically meaningful. Specifically, leveraging the interaction formulas from VinaDock, three types of edge with distance thresholds set to 2.7 Å, 3.4 Å, and 5 Å are established. The first two thresholds are intended to capture steric and short-range interactions, while the 5 Å threshold is specifically tailored to describe long-range interactions, such as van der Waals forces. These thresholds ensure that the model accurately captures different types of interactions based on spatial proximity, enhancing the precision of ligand generation conditioned on protein structures. We apply the same graph attention modules as in the global stage, but restrict attention to explicitly connected node pairs, enabling fine-grained interaction modeling between ligand atoms and their spatially proximal protein neighbors.
@@ 5.1 Experimental Setup
We use the CrossDocked [34] dataset, comprising 22.5 million docked protein–ligand complexes. Following standard protocols [35, 32, 28], we select samples with Root-Mean-Square Deviation (RMSD) << 1 Å and sequence identity << 30%. This yields 100,000 complexes for training and 100 for testing. For each test complex, 100 ligands are generated to ensure robust evaluation.
@@
Binding Affinity. Following prior works, we use AutoDock Vina [33] to evaluate binding energy from three perspectives: (a) Vina Score, which measures the energy of the original docked pose; (b) Vina Min, referring to the minimum energy after local minimization; and (c) Vina Dock, which represents the energy obtained through global redocking. Additionally, we compute Evina ($\downarrow$), the mean binding energy across all generated ligands; IMP% ($\uparrow$), the percentage of generated ligands outperforming the reference; and MPBG% ($\uparrow$), the average binding energy improvement over the reference.
@@
Structural Plausibility. Following Guan et al. [23], we compare the distributions of bond lengths, C–C distances within 2 Å, and all-atom distances within 12 Å. Bond length categories include C–C, C=C, C:C, C–N, C=N, C:N, C–O, and C=O, where the symbols “–”, “=”, and “:” respectively denote single, double, and aromatic bonds. These distributions are quantified using Jensen–Shannon Divergence, reported as JSD_BL ($\downarrow$), JSD_CC_2Å ($\downarrow$), and JSD_ALL_12Å ($\downarrow$), where lower values indicate better alignment with empirical statistics.
Conformational Stability. Following Harris et al. [36], we assess the conformational stability of generated ligands using two metrics: Strain Energy (SE) ($\downarrow$), which quantifies the internal energetic stability of the ligand and is reported at the 25th, 50th, and 75th percentiles, and Steric Clashes (Clash) ($\downarrow$), which measures the number of atomic overlaps between ligand and protein, reflecting spatial feasibility within the binding pocket.
@@ 6 Conclusion and Future Works
Suspendisse vel felis. Ut lorem lorem, interdum eu, tincidunt sit amet, laoreet vitae, arcu. Aenean faucibus pede eu ante. Praesent enim elit, rutrum at, molestie non, nonummy vel, nisl. Ut lectus eros, malesuada sit amet, fermentum eu, sodales cursus, magna. Donec eu purus. Quisque vehicula, urna sed ultricies auctor, pede lorem egestas dui, et convallis elit erat sed nulla. Donec luctus. Curabitur et nunc. Aliquam dolor odio, commodo pretium, ultricies non, pharetra in, velit. Integer arcu est, nonummy in, fermentum faucibus, egestas vel, odio.
@@ Acknowledgements
Hierarchical modeling has been extensively applied in molecular graph learning. HiMol[37] and HimGNN [38] partition molecules into atom, motif, and graph levels, and perform hierarchical aggregation for property prediction. HiGLLM [39] explores hierarchical graph representations in the context of large language models. ISA-PN [40] improves interpretability through positive and negative attention over hierarchical structures, while HMSN-CAM [41] employs hierarchical co-attention to enhance drug–drug interaction (DDI) prediction. However, these approaches operate solely within molecular graphs and rely on fixed, rule-based motif partitions. They are primarily designed for small molecules and are not well suited to the complex, multi-scale structural context encountered in protein-conditioned settings. In structure-based drug design (SBDD), protein structures involve more flexible and diverse hierarchies, including local residue clusters, functional regions, and long-range interactions that are difficult to capture with simple partitioning rules.
@@
Several prior works have leveraged protein surface features for learning molecular or protein representations. For example, [42] and [43] integrate surface geometry with sequence and structural information to obtain global protein embeddings, while [44] learns surface-based point cloud representations for protein-level prediction. However, these methods focus on representation learning and do not address molecular generation or protein–ligand interactions.
SurfGen [20] is more closely related to our setting, as it generates molecules based on protein surfaces. However, it models the protein solely at the surface level without incorporating atomic-level structural detail. Moreover, its autoregressive generation paradigm leads to slower inference and may increase the risk of generating redundant structures.
@@
*Autoregressive. Pocket2Mol [18] focuses on pocket features; GRAPHBP [17] employs graph-based ligand generation; and FLAG [19] uses fragment motifs.
*Non-autoregressive. We include diffusion-based models such as TARGETDIFF [23], DIFFBP [22], and DIFFSBDD [24]; D3FG [25], which integrates fragment motifs; DECOMPDIFF [26], employing arm and scaffold clustering; and MOLCRAFT [28], which is based on Bayesian Flow Networks. This comprehensive selection ensures a robust comparison across diverse generative methodologies.
@@
In training SculptDrug, we adopt the core BFNs configurations from MolCraft[28], utilizing 1,000 discrete time-steps to compute the discrete-time loss. The noise schedules are implemented with $\beta_{1}=1.5$ for atom types and $\sigma_{1}=0.03$ for atomic coordinate perturbations. The model is optimized over 26 epochs using Adam [45] with a learning rate of 0.005, momentum coefficients $(\beta_{1},\beta_{2})=(0.95,0.999)$, and a batch size of 32. Additionally, we apply an exponential moving average (EMA) across all parameters with a decay rate of 0.999. All experiments are executed on an NVIDIA A800 GPU (80GB VRAM) with CUDA 12.4 acceleration.
@@
To systematically evaluate the impact of hyperparameter choices on molecular generation performance, we set the number of clusters in the k-means++ [31] algorithm to one of the following four values: $N_{p}/4$, $N_{p}/8$, $N_{p}/16$, $N_{p}/32$, where $N_{p}$ denotes the total number of atoms in the protein binding pocket. Experimental results (Figure 5) show that setting $K=N_{p}/8$ achieves an optimal balance between global structural modeling and local interaction refinement. Specifically, when $K=N_{p}/4$, the JSD score increases significantly , suggesting that the large number of virtual atoms introduces excessive noise. In contrast, $K=N_{p}/32$ leads to a relatively poor Vina score and the highest clash score, indicating that the overly coarse granularity fails to capture crucial local structural details. Our experiments demonstrate that choosing $K=N_{p}/8$ preserves the essential local details while ensuring global structural consistency, thereby significantly enhancing ligand generation performance. Therefore, we adopt $K=N_{p}/8$ as the default configuration.

### v2

@@ 1 Introduction
Drug discovery is the process of identifying potential therapeutic molecules to treat diseases, playing a crucial role in improving health and addressing unmet medical needs [segler2018generating]. Yet, this process is highly complex and time-consuming. Structure-Based Drug Design (SBDD) has emerged as a powerful approach to streamline this process [anderson2003process], leveraging the structural information of biological targets to rationally design molecules with improved specificity and efficacy [isert2023structure].
The efficacy of SBDD lies in its ability to leverage the lock-and-key principle of protein-ligand interactions [jorgensen1991rusting], providing a framework for designing molecules that precisely target biological systems, as shown in Figure 1. Specifically, the protein surface functions as a unique “lock”, while the ligand serves as the complementary “key”. This structural compatibility ensures precise protein-ligand interactions, highlighting the effectiveness of SBDD in developing targeted and efficient drugs.
@@
Traditionally, SBDD relies on virtual screening, which aims to select ligands from an immense chemical space estimated to contain approximately $10^{60}$ potential molecules [reymond2012enumeration]. virtual screening focuses on predefined compound libraries based on molecular properties and structural features [lionta2014structure], which cover only a small fraction of the space and significantly limit ligand discovery [reymond2012enumeration]. Recent deep generative models [kingma2013auto, goodfellow2020generative, ho2020denoising, bfns] enable ligand generation beyond predefined libraries, supporting broader exploration of chemical space [cheng2021molecular].
@@
Diffusion models and Bayesian Flow Networks (BFNs) [bfns] are powerful tools for molecular generation, but often struggle with (3) Spatial Modeling Fidelity. Unlike images with fixed spatial layouts, molecules rely on flexible interatomic distances to define structure. Noise added to atomic coordinates can distort these distances, causing atoms to fall outside interaction thresholds and be excluded from local computations, ultimately compromising chemical validity.
@@ 2 Related Work
Ligand generation has evolved from 1D representations like SMILES [bjerrum2017molecular, segler2018generating], to 2D molecular graphs [liu2018constrained, jin2018junction], and more recently to 3D structure-based approaches. While 1D/2D methods capture some chemical constraints, they lack the spatial detail essential for modeling protein-ligand interactions. With the rise of deep learning in 3D modeling, structure-aware ligand generation has become a central research focus. Existing methods can be broadly categorized into: (1) Voxel-based and (2) Euclidean space-based models.
Voxel-based methods discretize 3D space into regular grids and model ligand density distributions within this voxelized space. Representative works such as LIGAN [ligan] and VoxBind [voxbind] generate ligands aligned to protein binding pockets using voxel-based representations. However, these methods suffer from high computational cost at fine resolutions and limited spatial precision due to the loss of structural detail during voxelization.
Euclidean space-based methods generate atomic types and positions directly in continuous 3D coordinates, typically representing proteins and ligands as point clouds or graphs. Early models, including GraphBP [Graphbp], Pocket2Mol [pocketmol], and FLAG [Flag], adopt autoregressive strategies at either the atom or fragment level. SurfGen [surfgen] and ResGen [resgen] incorporate protein context through surface features or residue-level graphs. However, the autoregressive nature of these methods can lead to error accumulation and limited global awareness.
Diffusion-based approaches, such as DIFFBP [Diffbp], TARGETDIFF [targetdiff], and DIFFSBDD [diffsbdd], leverage 3D-equivariant denoising to iteratively refine atom placements, improving spatial consistency and structural integrity. Extensions like D3FG [D3fg] and DECOMPDIFF [Decompdiff] introduce fragment priors and scaffold decomposition to enhance diversity and controllability. Nonetheless, modeling discrete atom types alongside continuous coordinates remains challenging. The introduction of Bayesian Flow Networks adds a new dimension to molecular generation  [GeoBFN]. MOLCRAFT [molcraft] generates ligands in continuous parameter spaces. However, their limited ability to perceive protein spatial geometry constrains performance in structure-guided ligand design.
@@ 4.1 Boundary Awareness Block
Inspired by SurfGen [surfgen] and SurfPro [song2024surfpro], we extract the surface structure of the binding pocket by selecting residues within 10 Å of any ligand atom. The solvent-excluded surface is computed using MSMS [msms], which outputs a triangular mesh of surface vertices and faces. We retain only the inward-facing vertices and their adjacent edges that are spatially close to the ligand. Each vertex is annotated with geometric and biochemical features: the shape index is computed from local curvature, while hydrophobicity, polarity, and electrostatic charge are assigned based on the nearest residue. The resulting surface graph aligns with the formal definition of $\mathcal{S}$.
@@ 4.2 Hierarchical Encoder
While atomic point clouds provide fine-grained local details, they often fail to capture higher-order structural patterns and are sensitive to noise. To address this, we first generate virtual atoms by clustering protein atoms, producing a coarse-grained representation that aggregates within-cluster features and highlights global structure. Ligands then interact with these virtual atoms to integrate global context, while fine-grained interactions are captured via multi-type edges that model detailed biochemical cues. To support both levels of interaction, we adopt equivariant attention mechanisms in the global and local modules, drawing inspiration from the attention design in DecompDiff [Decompdiff].
@@
To obtain a coarse-grained protein representation, we apply k-means++ [arthur2006k] clustering to protein atoms based on 3D coordinates, generating virtual atoms at cluster centroids. Let $C$ denote the atoms in a cluster. The virtual atom’s position is defined as:
@@
Note: Evina >> 0 are considered invalid and are represented by “–”. Baseline ligand results are from CBGbench [Cbgbench].
@@
We designed fine-grained edge connection rules grounded in established domain knowledge, specifically inspired by the affinity calculations in AutoDock Vina [AutoDock]. This approach ensures that our modeling of molecular interactions is both accurate and biologically meaningful. Specifically, leveraging the interaction formulas from VinaDock, three types of edge with distance thresholds set to 2.7 Å, 3.4 Å, and 5 Å are established. The first two thresholds are intended to capture steric and short-range interactions, while the 5 Å threshold is specifically tailored to describe long-range interactions, such as van der Waals forces. These thresholds ensure that the model accurately captures different types of interactions based on spatial proximity, enhancing the precision of ligand generation conditioned on protein structures. We apply the same graph attention modules as in the global stage, but restrict attention to explicitly connected node pairs, enabling fine-grained interaction modeling between ligand atoms and their spatially proximal protein neighbors.
@@ 5.1 Experimental Setup
We use the CrossDocked [CrossDocked] dataset, comprising 22.5 million docked protein–ligand complexes. Following standard protocols [luo20213d, Cbgbench, molcraft], we select samples with Root-Mean-Square Deviation (RMSD) << 1 Å and sequence identity << 30%. This yields 100,000 complexes for training and 100 for testing. For each test complex, 100 ligands are generated to ensure robust evaluation.
@@
Binding Affinity. Following prior works, we use AutoDock Vina [AutoDock] to evaluate binding energy from three perspectives: (a) Vina Score, which measures the energy of the original docked pose; (b) Vina Min, referring to the minimum energy after local minimization; and (c) Vina Dock, which represents the energy obtained through global redocking. Additionally, we compute Evina ($\downarrow$), the mean binding energy across all generated ligands; IMP% ($\uparrow$), the percentage of generated ligands outperforming the reference; and MPBG% ($\uparrow$), the average binding energy improvement over the reference.
@@
Structural Plausibility. Following Guan et al. [targetdiff], we compare the distributions of bond lengths, C–C distances within 2 Å, and all-atom distances within 12 Å. Bond length categories include C–C, C=C, C:C, C–N, C=N, C:N, C–O, and C=O, where the symbols “–”, “=”, and “:” respectively denote single, double, and aromatic bonds. These distributions are quantified using Jensen–Shannon Divergence, reported as JSD_BL ($\downarrow$), JSD_CC_2Å ($\downarrow$), and JSD_ALL_12Å ($\downarrow$), where lower values indicate better alignment with empirical statistics.
Conformational Stability. Following Harris et al. [posecheck], we assess the conformational stability of generated ligands using two metrics: Strain Energy (SE) ($\downarrow$), which quantifies the internal energetic stability of the ligand and is reported at the 25th, 50th, and 75th percentiles, and Steric Clashes (Clash) ($\downarrow$), which measures the number of atomic overlaps between ligand and protein, reflecting spatial feasibility within the binding pocket.
@@ Acknowledgements
Hierarchical modeling has been extensively applied in molecular graph learning. HiMol[HiMol] and HimGNN [himgnn] partition molecules into atom, motif, and graph levels, and perform hierarchical aggregation for property prediction. HiGLLM [HiGLLM] explores hierarchical graph representations in the context of large language models. ISA-PN [ISA-PN] improves interpretability through positive and negative attention over hierarchical structures, while HMSN-CAM [HMSN-CAM] employs hierarchical co-attention to enhance drug–drug interaction (DDI) prediction. However, these approaches operate solely within molecular graphs and rely on fixed, rule-based motif partitions. They are primarily designed for small molecules and are not well suited to the complex, multi-scale structural context encountered in protein-conditioned settings. In structure-based drug design (SBDD), protein structures involve more flexible and diverse hierarchies, including local residue clusters, functional regions, and long-range interactions that are difficult to capture with simple partitioning rules.
@@
Several prior works have leveraged protein surface features for learning molecular or protein representations. For example, [pretrain_surface] and [surface_vqmae] integrate surface geometry with sequence and structural information to obtain global protein embeddings, while [surface_cvpr] learns surface-based point cloud representations for protein-level prediction. However, these methods focus on representation learning and do not address molecular generation or protein–ligand interactions.
SurfGen [surfgen] is more closely related to our setting, as it generates molecules based on protein surfaces. However, it models the protein solely at the surface level without incorporating atomic-level structural detail. Moreover, its autoregressive generation paradigm leads to slower inference and may increase the risk of generating redundant structures.
@@
*Autoregressive. Pocket2Mol [pocketmol] focuses on pocket features; GRAPHBP [Graphbp] employs graph-based ligand generation; and FLAG [Flag] uses fragment motifs.
*Non-autoregressive. We include diffusion-based models such as TARGETDIFF [targetdiff], DIFFBP [Diffbp], and DIFFSBDD [diffsbdd]; D3FG [D3fg], which integrates fragment motifs; DECOMPDIFF [Decompdiff], employing arm and scaffold clustering; and MOLCRAFT [molcraft], which is based on Bayesian Flow Networks. This comprehensive selection ensures a robust comparison across diverse generative methodologies.
@@
In training SculptDrug, we adopt the core BFNs configurations from MolCraft[molcraft], utilizing 1,000 discrete time-steps to compute the discrete-time loss. The noise schedules are implemented with $\beta_{1}=1.5$ for atom types and $\sigma_{1}=0.03$ for atomic coordinate perturbations. The model is optimized over 26 epochs using Adam [Kingma2014AdamAM] with a learning rate of 0.005, momentum coefficients $(\beta_{1},\beta_{2})=(0.95,0.999)$, and a batch size of 32. Additionally, we apply an exponential moving average (EMA) across all parameters with a decay rate of 0.999. All experiments are executed on an NVIDIA A800 GPU (80GB VRAM) with CUDA 12.4 acceleration.
@@
To systematically evaluate the impact of hyperparameter choices on molecular generation performance, we set the number of clusters in the k-means++ [arthur2006k] algorithm to one of the following four values: $N_{p}/4$, $N_{p}/8$, $N_{p}/16$, $N_{p}/32$, where $N_{p}$ denotes the total number of atoms in the protein binding pocket. Experimental results (Figure 5) show that setting $K=N_{p}/8$ achieves an optimal balance between global structural modeling and local interaction refinement. Specifically, when $K=N_{p}/4$, the JSD score increases significantly , suggesting that the large number of virtual atoms introduces excessive noise. In contrast, $K=N_{p}/32$ leads to a relatively poor Vina score and the highest clash score, indicating that the overly coarse granularity fails to capture crucial local structural details. Our experiments demonstrate that choosing $K=N_{p}/8$ preserves the essential local details while ensuring global structural consistency, thereby significantly enhancing ligand generation performance. Therefore, we adopt $K=N_{p}/8$ as the default configuration.